OpenStudy (anonymous):

Not sure how to solve this? Any help? :) Sharmi and Reese each deposit money into a savings account. Sharmi deposits $300 less than Reese. After the same amount of time, Reese has earned $180 in interest at a rate of 3% per year, and Sharmi has earned $270 in interest at a rate of 6% per year. How much money did Reese and Sharmi initially deposit into their savings accounts?

ganeshie8 (ganeshie8):

lets say, reese deposts : x sharmi deposited : x-300

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

reese earned $180 interest on x deposit at 3% rate lets say number of years = t, apply Simple Interest forumula : 180 = P*R*T 180 = x*.03*t ----------------------(1)

ganeshie8 (ganeshie8):

sharmi earned $270 interest on (x-300) deposit at 6% rate since time is same, number of years = t, apply Simple Interest forumula : 270 = P*R*T 270 = (x-300)*.03*t ----------------------(2)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

you have two equations, two unknowns, x & t. you can solve them for x ?

ganeshie8 (ganeshie8):

oops ! second equation has a typo. it should be : 270 = (x-300)*\(\color{red}{.06}\)*t ----------------------(2)

OpenStudy (anonymous):

i'm a little bit confused, could you show me what the equations look like?

ganeshie8 (ganeshie8):

from (1) t = 180/x*.03 substitue this value of t in (2) and solve for x

ganeshie8 (ganeshie8):

270 = (x-300) * .06 * \(\frac{180}{x*.03}\)

ganeshie8 (ganeshie8):

270x = (x-300) *2 * 180

OpenStudy (anonymous):

so it looks like t=180/2(.03) ?

ganeshie8 (ganeshie8):

hmm how did u get this ?

OpenStudy (anonymous):

i'm not sure exactly